The density function of X is given by f(x) = a + bx2 if 0 ? x ? 1 0 otherwise. Suppose also that you are told that E(X) = 3/5. (a) Find a and b. (b) Determine the cdf, F(x), explicilty.Problem 4. The density function of X is given by f(z) = 0 otherwise. Suppose also that you are told that E(X-3/5. (a) Find a and b. b) Determine the cdf, F(r
Answer #1

According to the given question:

x is a continious variable with the following probability density function as:

f(x)=a+bx^{2} if  0\leq x\leq 1

=0 otherwise.

and given expectation is

E(x)=\frac{3}{5}

i) According to the question, from the property of density function as:

1)\int_{-\alpha }^{\alpha }f(x)dx=1

2)f(x)\geq 0 for all x

Therefore from the property of density function:

\int_{-\alpha }^{\alpha }(a+bx^{2})dx=1

or, \int_{0 }^{1 }(a+bx^{2})dx=1

or, \left |ax+\frac{bx^{2+1}}{3} \right |_{0}^{1}=1

or, (a+\frac{b}{3})=1....(i)

from the expectation we get,

E(x)=\int_{-\alpha }^{\alpha }x.(a+bx^{2})dx=\frac{3}{5}

or, \int_{0 }^{1 }x(a+bx^{2})dx=\frac{3}{5}

or, \left |a\frac{x^{2}}{2}+\frac{bx^{3+1}}{4} \right |_{0}^{1}=\frac{3}{5}

or, (\frac{a}{2}+\frac{b}{4})=\frac{3}{5}....(ii)

from equation (i) and (ii) we get, from (ii)*2 we get,

(a+\frac{b}{2})=\frac{6}{5}...(iii)

now (i) -(ii) we get;

(\frac{b}{3}-\frac{b}{2})=1-\frac{6}{5}

or, \frac{-b}{6}=-\frac{1}{5}

or, b=\frac{6}{5}

Therefore a is determine as:

a+\frac{b}{3}=1

or, a+\frac{2}{5}=1

or, a=1-\frac{2}{5}

or a=\frac{3}{5}

Therefore the probability density fucntion is :

f(x)=\frac{3}{5}+\frac{6}{5}x^{2} if  0\leq x\leq 1

=0 otherwise.

b) Now the cumulative density function (CDF) is obtained by integrating the probability density function as shown below:

F(x)=\int_{-\alpha }^{x }f(t)dt

  =\int_{0 }^{x}(\frac{3}{5}+\frac{6}{5}t^{2})dt

  =\left |\frac{3t}{5}+\frac{6t^{3}}{15} \right |_{0}^{x}

Thereore,

F(x)=(\frac{3x}{5}+\frac{2x^{3}}{5})

Therefore the required CDF is as follows:

F(x)=0 when x<0

=(\frac{3x}{5}+\frac{2x^{3}}{5}) when  0\leq x\leq 1

=1 when   x\geq 1